# Particle Filtering for a Class of State-Space Models with Low and Degenerate Observational Noise

### v1

@@ 3.1.1 Low Noise Case
In such a scenario, one can use a standard particle filter, which we now detail. Let $\widetilde{q}_{n}^{\Delta}$ be a sequence of positive Markov transition densities on $\mathbb{R}^{d_{x}}$ and set for any $(z^{\prime},z)\in\mathbb{R}^{2d_{x}}$
$w_{n}^{\Delta}(z^{\prime},z)=\frac{p_{n}(u_{n}^{\Delta}(z,\epsilon))f_{n}(u_{n}^{\Delta}(z,x)|u_{n-1}^{\Delta}(z^{\prime},x))}{\widetilde{q}_{n}^{\Delta}(z|z^{\prime})}.$
Note that $\widetilde{q}_{n}^{\Delta}$ can depend on $\Delta$ and when $n=1$ we may (or may not) condition on $x_{0}$ in $\widetilde{q}_{n}^{\Delta}$ but write $\widetilde{q}_{n}^{\Delta}(\cdot|x_{0})$ from herein. Let $N\in\mathbb{N}$ be given and denote for any $n\geq 1$, $\widetilde{z}_{n}^{1:N}=(\widetilde{z}_{n}^{1}\dots,\widetilde{z}_{n}^{N})$ which represent the particles of the PF. It will prove useful to use the ancestor representation of the resampling indices so we will also be interested in $(\widetilde{a}_{n}^{1},\dots,\widetilde{a}_{n}^{N})\in\{1,\dots,N\}^{N}$ and we use the convention that $\widetilde{a}_{0}^{i}=i$ for each $i\in\{1,\dots,N\}$. We will also write $\widetilde{z}_{0}^{\widetilde{a}_{0}^{j}}=x_{0}$. The PF is then the following.
  • Initialize: For $i\in\{1,\dots,N\}$, sample $\widetilde{Z}_{1}^{i}|x_{0}$ i.i.d. from $\widetilde{q}_{1}^{\Delta}(\cdot|x_{0})$. Set $n=2$.
  • Iterate: For $i\in\{1,\dots,N\}$, sample $\widetilde{a}_{n-1}^{i}|\widetilde{z}_{n-2}^{1:N},\widetilde{z}_{n-1}^{1:N}$ i.i.d. from
    $\frac{w_{n-1}^{\Delta}(\widetilde{z}_{n-2}^{\widetilde{a}_{n-2}^{i}},\widetilde{z}_{n-1}^{\widetilde{a}_{n-1}^{i}})}{\sum_{s=1}^{N}w_{n-1}^{\Delta}(\widetilde{z}_{n-2}^{\widetilde{a}_{n-2}^{s}},\widetilde{z}_{n-1}^{s})}.$
@@
This can only be used in certain scenarios, one of which we investigate below.
3.1.2 Degenerate Noise Case
@@
where $V_{n}$ is the $d_{x}\times(d_{x}-d_{y})$ matrix whose columns form an orthonormal basis of $\text{ker}(A_{n})=\{x\in\mathbb{R}_{d_{x}}:A_{n}x=0\}$ and with the convention that $u_{0}^{\star}(z)=x_{0}$ for any $z\in\mathbb{R}^{d_{x}-d_{y}}$. At any time $n\geq 1$ we can write the joint smoothing density of $z_{1:n}\in\mathbb{R}^{n(d_{x}-d_{y})}$ w.r.t. $\bigotimes_{i=1}^{n}\lambda_{d_{x}-d_{y}}(dz_{i})$ as
$\pi_{n}^{\star}(z_{1:n})\propto\prod_{k=1}^{n}f_{k}(u_{k}^{\star}(z_{k})|u_{k-1}^{\star}(z_{k-1}))$
@@
  • Iterate: For $i\in\{1,\dots,N\}$, sample $a_{n-1}^{i}|z_{n-2}^{1:N},z_{n-1}^{1:N}$ i.i.d. from
    $\frac{w_{n-1}^{\star}(z_{n-2}^{a_{n-2}^{i}},z_{n-1}^{a_{n-1}^{i}})}{\sum_{s=1}^{N}w_{n-1}^{\star}(z_{n-2}^{a_{n-2}^{s}},z_{n-1}^{s})}.$
@@
3.1.3 Mathematical Result
We now consider the convergence, in a sense to be made precise below, of the PF in the low noise case to that of the degenerate noise case. Recall that we use the notation for $\widetilde{z}\in\mathbb{R}^{d_{x}}$ as $\widetilde{z}^{\top}=(z^{\top},\overline{z}^{\top})$, with $z\in\mathbb{R}^{d_{x}}$. We make the following assumption.
@@ 3.1.3 Mathematical Result
This assumption is not overly restrictive as one chooses the proposal densities. The assumption means that as the noise disappears, one should propose the $d_{x}-d_{y}$ random variable, just as in the in the degenerate noise case and the noise variable according to the noise density.
@@
The main purpose of Proposition 3.1 is to verify that the properties of the low noise case PF are inherited, as $\Delta\downarrow 0$ to that of the degenerate noise case. In other words if one can design a good PF for the degenerate noise case, then the low noise filtering problem can be effectively handled. For instance if one has that
@@
$\sup_{n\geq 1}\mathbb{E}^{\star}\left[\left|\pi_{n}^{\star,N}(\varphi)-\pi_{n}^{\star}(\varphi)\right|^{r}\right]^{1/r}=\mathcal{O}(N^{-\tfrac{1}{2}})$
and so the time uniform errors of the degenerate noise PF are then then obtained for the low noise PF.
3.2 Special Case of Additive Noise
@@
$X_{n}$ $=f_{n}(X_{n-1})+\nu_{n}$
where $Y_{n}\in\mathbb{R}^{d_{y}}$, $1\leq d_{y}<d_{x}$, $A_{n}\in\mathbb{R}^{d_{y}\times d_{x}}$ is of full row rank, $\Delta\geq 0$ $f_{n}:\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}^{d_{x}}$. Independently for each $n\in\mathbb{N}$ $\epsilon_{n}\sim\mathcal{N}_{d_{y}}(0,\Sigma_{n})$ and independently $\nu_{n}\sim\mathcal{N}_{d_{x}}(0,\Omega_{n})$ where, $\Sigma_{n},\Omega_{n}$ are a sequence of real positive definite symmetric matrices and for $d\in\mathbb{N}$ fixed, $(\mu,\kappa)\in\mathbb{R}^{d}\times\mathbb{R}^{d\times d}$ ($\kappa$ is also positive definite and symmetric) $\mathcal{N}_{d}(\mu,\kappa)$ is the $d$-dimensional Gaussian distribution with mean $\mu$ and covariance $\kappa$.
3.2.1 Particle Filter for the Natural Parameterization
For this problem we already know the PF for the filter in the noise parameterization case as well as the degenerate noise case. For completeness, we detail the natural parameterization case when $\Delta>0$. $q_{n}$ is a sequence of positive Markov transition densities on $\mathbb{R}^{d_{x}}$ and we set for any $(x^{\prime},x)\in\mathbb{R}^{2d_{x}}$
As before we set $a_{0}^{i}=i$ for $i\in\{1,\dots,n\}$ and $x_{0}^{i}=x_{0}$. The PF is then the following.
  • Initialize: For $i\in\{1,\dots,N\}$, sample $X_{1}^{i}|x_{0}$ i.i.d. from $q_{1}(\cdot|x_{0})$. Set $n=2$.
  • Iterate: For $i\in\{1,\dots,N\}$, sample $a_{n-1}^{i}|x_{n-2}^{1:N},x_{n-1}^{1:N}$ i.i.d. from
    $\frac{w_{n-1}(x_{n-2}^{a_{n-2}^{i}},x_{n-1}^{a_{n-1}^{i}})}{\sum_{s=1}^{N}w_{n-1}(x_{n-2}^{a_{n-2}^{s}},x_{n-1}^{s})}.$
The optimal proposal in this case is
3.2.2 Mathematical Result
We now consider the optimal or bootstap proposals and the importance weights in the natural parameterization, noise parameterization (low noise) and the degenerate noise case. To aid the notations in the below statement we partition the matrix $V_{n}^{\Delta}$ into its first $d_{x}$ rows denoted by the $d_{x}\times d_{x}$ matrix $V_{n}^{\Delta}(x)$ and its last $d_{y}$ rows denoted by the $d_{y}\times d_{x}$ matrix $V_{n}^{\Delta}(\epsilon)$. We will also use the notation $x_{0}^{\star}+V^{\Delta}_{0}(x)z=x_{0}$ $x_{0}^{\star}+V_{0}z=x_{0}$ and where we have not defined the objects on L.H.S.. We have the following result whose proof is in Appendix A.2.
@@ 3.2.2 Mathematical Result
    $w_{n}(x_{n})$ $=\exp\left\{-\frac{1}{2\Delta}(y_{n}-A_{n}x_{n})^{\top}\Sigma^{-1}(y_{n}-A_{n}x_{n})\right\}$
@@
The significance of Proposition 3.2 is as follows. $\Delta$ controls the amount of noise and the low noise case is as $\Delta\downarrow 0$. In the case of the bootstrap PF it is clear that numerically that $w_{n}(x_{n})$ will degenerate to zero and hence this algorithm is of little use; this is unsuprising as the proposal does not take into account any information associated to the data. We remark that for the optimal proposal in the case of the natural parameterization the weight does not collapse or explode, but clearly the proposal and target (which begins to concentrate on $\mathsf{M}_{n}$) will become mutually singular as $\Delta\downarrow 0$ so it is of little use in the very small noise and degenerate case. In the low and degenerate noise cases (2 & 3 in Proposition 3.2) one can characterize the optimal proposal and the associated weight and, as was the case in Proposition 3.1, there is a convergence of the weights from the low noise to degenerate case with a similar conclusion to Remark 3.2.
@@ 4.1 Diffusion Process
where $X_{t}\in\mathbb{R}^{d_{x}}$, $\mu:\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}^{d_{x}}$, and $\sigma:\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}^{d_{x}\times d_{x}}$ and $\{W_{t}\}_{t\geq 0}$ is a standard $d_{x}$-dimensional Brownian motion. Denote $\mu^{j}$ as the $j^{th}$-component of $\mu$, $j\in\{1,\dots,d_{x}\}$ and $\sigma^{j,k}$ as the $(j,k)^{th}$-component of $\sigma$, $(j,k)\in\{1,\dots,d_{x}\}^{2}$. As the exposition will closely follow the works in [1, 4, 13, 17] we make the following assumption which is also made in most of those papers. Below $\mathcal{C}(\mathbb{R}^{d_{x}})$ are the twice continuously differentiable functions on $\mathbb{R}^{d_{x}}$ and $\|\cdot\|$ is the $L_{2}$-norm.
@@
We note that one could use the Euler-Maruyama time-discretization method to obtain the transition density of the hidden process. However, it is well-known (e.g. [4] and the references therein) that this method, when combined with a PF, typically collapses (in some sense) as the time-discretization becomes finer and thus we consider a bridge construction as often used in the literature. As a result the PFs that we will present in Sections 4.3.1 and 4.3.2 will be robust to the case of low noise and fine time discretizations.
@@ 4.2 Bridge Construction
We review the method in [16, 18] as has been used in [4]. The exposition closely follows that in [1, 4, 13]. Consider the case $t\in[s_{1},s_{2}]$, $0\leq s_{1},s_{2}\leq T$ and let $\mathbf{X}_{[s_{1},s_{2}]}:=\{X_{t}\}_{t\in[s_{1},s_{2}]}$, and $\mathbf{W}_{[s_{1},s_{2}]}:=\{W_{t}\}_{t\in[s_{1},s_{2}]}$. Set $f_{t,s_{2}}(x^{\prime}|x)$ denote the unknown transition density from time $t$ to $s_{2}$ associated to (4.1). If one could sample from $f_{s_{1},s_{2}}$ to obtain $(x,x^{\prime})\in\mathbb{R}^{2d_{x}}$, we will explain that we can interpolate these points by using a bridge process. This process will have a drift given by $\mu(x)+\Sigma(x)\nabla_{x}\log{f}_{t,s_{2}}(x^{\prime}|x)$. Let $\mathbb{P}_{x,x^{\prime}}$ denote the law of the solution of the SDE (4.1), on $s_{1},s_{2}]$, started at $x$ and conditioned to hit $x^{\prime}$ at time $s_{2}$.
@@
(4.2) is specified so that its transition density $\tilde{f}$ is available; see [16, Section 2.2] for the technical conditions on $\tilde{\mu},\tilde{\Sigma},\tilde{f}$. The main purpose of $\{\tilde{X}_{t}\}_{t\in[s_{1},s_{2}]}$ is to sample $x^{\prime}$ and use its transition density to construct another process $\{X_{t}^{\circ}\}_{t\in[s_{1},s_{2}]}$ conditioned to hit $x^{\prime}$ at $t=s_{2}$; which in turn will be an importance proposal for $\{X_{t}\}_{t\in[s_{1},s_{2}]}$. Set:
@@
[16] prove that $\mathbb{P}_{x,x^{\prime}}$ and $\mathbb{P}^{\circ}_{x,x^{\prime}}$ are absolutely continuous w.r.t. each other, with Radon-Nikodym derivative:
$\frac{d\mathbb{P}_{x,x^{\prime}}}{d\mathbb{P}^{\circ}_{x,x^{\prime}}}(\mathbf{X}_{[s_{1},s_{2}]})=\exp\Big\{\int_{s_{1}}^{s_{2}}L_{s_{2}}(t,X_{t})dt\Big\}\times\frac{\tilde{f}_{s_{1},s_{2}}(x^{\prime}|x)}{f_{s_{1},s_{2}}(x^{\prime}|x)},$ (4.5)
@@
with $\textrm{Tr}(\cdot)$ denoting the trace of a squared matrix. Set
$R_{s_{1},s_{2}}(\mathbf{X}_{[s_{1},s_{2}]}):=\frac{d\mathbb{P}_{x,x^{\prime}}}{d\mathbb{P}^{\circ}_{x,x^{\prime}}}(\mathbf{X}_{[s_{1},s_{2}]})f_{s_{1},s_{2}}(x^{\prime}|x).$

### v2

@@ 3.1.1 The Low Noise Case
In such a scenario, one can use a standard particle filter, which we now detail. Let $\widetilde{q}_{n}^{\Delta}$ be a sequence of positive Markov transition densities on $\mathbb{R}^{d_{x}}$ and set for any $(\widetilde{z}^{\prime},\widetilde{z})\in\mathbb{R}^{2d_{x}}$
$w_{n}^{\Delta}(\widetilde{z}^{\prime},\widetilde{z})=\frac{p_{n}(u_{n}^{\Delta}(\widetilde{z},\epsilon))f_{n}(u_{n}^{\Delta}(\widetilde{z},x)|u_{n-1}^{\Delta}(z^{\prime},x))}{\widetilde{q}_{n}^{\Delta}(\widetilde{z}|\widetilde{z}^{\prime})}$.
As for the natural parameterization case, note that $\widetilde{q}_{n}^{\Delta}$ can depend on $\Delta$ and when $n=1$ we may (or may not) condition on $x_{0}$ in $\widetilde{q}_{1}^{\Delta}$ but write $\widetilde{q}_{1}^{\Delta}(\cdot|x_{0})$ from herein. Let $N\in\mathbb{N}$ be given and denote for any $n\geq 1$, $\widetilde{z}_{n}^{1:N}=(\widetilde{z}_{n}^{1}\dots,\widetilde{z}_{n}^{N})$ which represent the particles of the PF. It will prove useful to use the ancestor representation of the resampling indices so we will also be interested in $(\widetilde{a}_{n}^{1},\dots,\widetilde{a}_{n}^{N})\in\{1,\dots,N\}^{N}$ and again we use the convention that $\widetilde{a}_{0}^{i}=i$ for each $i\in\{1,\dots,N\}$. We will also write $\widetilde{z}_{0}^{\widetilde{a}_{0}^{j}}=x_{0}$.
The PF is then the following algorithm:
  • Initialize: For $i\in\{1,\dots,N\}$, sample $\widetilde{Z}_{1}^{i}|x_{0}$ from $\widetilde{q}_{1}^{\Delta}(\cdot|x_{0})$. Set $n=2$.
  • Iterate: For $i\in\{1,\dots,N\}$, sample $\widetilde{a}_{n-1}^{i}\sim\text{Cat}(\cdot;N,\mathbf{w}^{\Delta}_{n-1})$ with
    $\mathbf{w}^{\Delta}_{n-1}=\left(\frac{w_{n-1}^{\Delta}(\widetilde{z}_{n-2}^{\widetilde{a}_{n-2}^{1}},\widetilde{z}_{n-1}^{1})}{\sum_{s=1}^{N}w_{n-1}^{\Delta}(\widetilde{z}_{n-2}^{\widetilde{a}_{n-2}^{s}},\widetilde{z}_{n-1}^{s})},\ldots,\frac{w_{n-1}^{\Delta}(\widetilde{z}_{n-2}^{\widetilde{a}_{n-2}^{N}},\widetilde{z}_{n-1}^{N})}{\sum_{s=1}^{N}w_{n-1}^{\Delta}(\widetilde{z}_{n-2}^{\widetilde{a}_{n-2}^{s}},\widetilde{z}_{n-1}^{s})}\right).$
@@
This is tractable only in certain modelling scenarios, one of which we investigate below. We remark that in the case that one is able to compute the optimal proposal, due to the one-to-one mappings from the natural parameterization to the noise parameterization to the map back to the Lebesgue measure in the linear constraint scenario, the optimal proposal from the natural parameterization will provide an equivalent formulation. One of the main points of the low noise method however, is that it will permit its application in cases where the optimal proposal is not available and will be robust to low values of the noise parameter.
3.1.2 The Degenerate Noise Case
@@
with the convention that $u_{0}^{\star}(z)=x_{0}$ for any $z\in\mathbb{R}^{d_{x}-d_{y}}$. At any time $n\geq 1$ we can write the joint smoothing density of $z_{1:n}\in\mathbb{R}^{n(d_{x}-d_{y})}$ w.r.t. $\bigotimes_{i=1}^{n}\lambda_{d_{x}-d_{y}}(dz_{i})$ as
$\pi_{n}^{\star}(z_{1:n})\propto\prod_{k=1}^{n}f_{k}(u_{k}^{\star}(z_{k})|u_{k-1}^{\star}(z_{k-1})),$
@@
  • Iterate: For $i\in\{1,\dots,N\}$, sample $a_{n-1}^{i}\sim\text{Cat}(\cdot;N,\mathbf{w}^{\star}_{n-1})$ with
    $\mathbf{w}^{\star}_{n-1}=\left(\frac{w_{n-1}^{\star}(z_{n-2}^{a_{n-2}^{1}},z_{n-1}^{1})}{\sum_{s=1}^{N}w_{n-1}^{\star}(z_{n-2}^{a_{n-2}^{s}},z_{n-1}^{s})},\ldots,\frac{w_{n-1}^{\star}(z_{n-2}^{a_{n-2}^{N}},z_{n-1}^{N})}{\sum_{s=1}^{N}w_{n-1}^{\star}(z_{n-2}^{a_{n-2}^{s}},z_{n-1}^{s})}\right).$
@@
3.1.3 A Low Noise Convergence Result
We now consider the convergence, in a sense to be made precise below, of the PF in the low noise case to that of the degenerate noise case. Recall that we use the notation for $\widetilde{z}\in\mathbb{R}^{d_{x}}$ as $\widetilde{z}^{\top}=(z^{\top},\overline{z}^{\top})$, with $z\in\mathbb{R}^{d_{x}-d_{y}}$. We make the following assumption.
@@ 3.1.3 A Low Noise Convergence Result
This assumption is not overly restrictive as one chooses the proposal densities. The assumption means that as the noise disappears, one should propose the $d_{x}-d_{y}$ random variable, just as in the degenerate noise case and the noise variable according to the noise density.
@@
The main purpose of Proposition 3.1 is to verify that the properties of the low noise case PF are inherited, as $\Delta\downarrow 0$ from that of the degenerate noise case. In other words if one can design a good PF for the degenerate noise case, then the low noise filtering problem can be effectively handled. For instance if one has that
@@
$\sup_{n\geq 1}\mathbb{E}^{\star}\left[\left|\pi_{n}^{\star,N}(\varphi)-\pi_{n}^{\star}(\varphi)\right|^{r}\right]^{1/r}\leq\frac{C}{N^{-\tfrac{1}{2}}}$
where $C<+\infty$ is a finite constant. Then for any fixed $n$ we would have
$\lim_{\Delta\downarrow 0}\mathbb{E}^{\Delta}\left[\left|\pi_{n}^{\Delta,N}(\varphi)-\pi_{n}^{\Delta}(\varphi)\right|^{r}\right]^{1/r}=\mathbb{E}^{\star}\left[\left|\pi_{n}^{\star,N}(\varphi)-\pi_{n}^{\star}(\varphi)\right|^{r}\right]^{1/r}\leq\frac{C}{N^{-\tfrac{1}{2}}}.$
3.2 Additive Gaussian Noise with Nonlinear Signal Dynamics
@@
$X_{n}$ $=f_{n}(X_{n-1})+\nu_{n},$
where $Y_{n}\in\mathbb{R}^{d_{y}}$, $1\leq d_{y}<d_{x}$, $A_{n}\in\mathbb{R}^{d_{y}\times d_{x}}$ is of full row rank, $\Delta\geq 0$ $f_{n}:\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}^{d_{x}}$. Independently for each $n\in\mathbb{N}$ $\epsilon_{n}\sim\mathcal{N}_{d_{y}}(0,\Sigma_{n})$ and independently $\nu_{n}\sim\mathcal{N}_{d_{x}}(0,\Omega_{n})$ where, $\Sigma_{n},\Omega_{n}$ are a sequence of real positive definite symmetric matrices. Here $\mathcal{N}_{d}(\mathsf{m},\mathsf{K})$ denotes the $d$-dimensional Gaussian distribution with mean $\mathsf{m}$ and covariance matrix $\mathsf{K}$.
3.2.1 Optimal Proposals
We now consider the optimal or bootstap proposals and the importance weights in the natural parameterization, noise parameterization (low noise) and the degenerate noise case. To aid the notations in the below statement we partition the matrix $V_{n}^{\Delta}$ into its first $d_{x}$ rows denoted by the $d_{x}\times d_{x}$ matrix $V_{n}^{\Delta}(x)$ and its last $d_{y}$ rows denoted by the $d_{y}\times d_{x}$ matrix $V_{n}^{\Delta}(\epsilon)$. We will also use the notation $x_{0}^{\star}+V^{\Delta}_{0}(x)z=x_{0}$ $x_{0}^{\star}+V_{0}z=x_{0}$ and where we have not defined the objects on left hand side. We have the following result whose proof is in Appendix A.3.
@@ 3.2.1 Optimal Proposals
    $w_{n}(x_{n})$ $=\exp\left\{-\frac{1}{2\Delta}(y_{n}-A_{n}x_{n})^{\top}\Sigma_{n}^{-1}(y_{n}-A_{n}x_{n})\right\}$
@@
The significance of Proposition 3.2 is as follows. $\Delta$ controls the amount of noise and the low noise case is as $\Delta\downarrow 0$. In the case of the bootstrap PF it is clear that numerically that $w_{n}(x_{n})$ will degenerate to zero and hence this algorithm is of little use; this is unsurprising as the proposal does not take into account any information associated to the data. In the low and degenerate noise cases (2 & 3 in Proposition 3.2) one can characterize the optimal proposal and the associated weight and, as was the case in Proposition 3.1, there is a convergence of the weights from the low noise to degenerate case with a similar conclusion to Remark 3.2. As we noted before the optimal proposal in the natural parameterization will be equivalent to the one in the noise parameterization.
@@ 4.1 Diffusion Process
where $X_{t}\in\mathbb{R}^{d_{x}}$, $\mu:\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}^{d_{x}}$, and $\sigma:\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}^{d_{x}\times d_{x}}$ and $\{W_{t}\}_{t\geq 0}$ is a standard $d_{x}$-dimensional Brownian motion. Denote $\mu^{j}$ as the $j^{th}$-component of $\mu$, $j\in\{1,\dots,d_{x}\}$ and $\sigma^{j,k}$ as the $(j,k)^{th}$-component of $\sigma$, $(j,k)\in\{1,\dots,d_{x}\}^{2}$. As the exposition will closely follow the works in [1, 4, 15, 25] we make the following assumption which is also made in most of those papers. Below $\mathcal{C}(\mathbb{R}^{d_{x}})$ are the twice continuously differentiable functions on $\mathbb{R}^{d_{x}}$ and $\|\cdot\|$ is the Euclidean norm.
@@
We note that one could use the Euler-Maruyama time-discretization method to obtain the transition density of the hidden process. However, it is well-known (e.g. [4] or [27, Figure 1]) that this method, when combined with a PF, typically collapses (due to increasing Monte Carlo variance) as the time-discretization becomes finer and thus we consider a robust reformulation using a bridge construction that avoids this issue and is often used in the literature [24, 26]. Both approaches will need to be discretized in time, but will opt for the latter due to its robustness to fine time discretizations. As a result the PFs that we will present in Sections 4.3.1 and 4.3.2 will be robust both to the case of low observation noise and fine time discretizations.
@@ 4.2 Bridge Construction
We review the method in [24, 26] as has been used in [4]. The exposition closely follows that in [1, 4, 15]. Consider the case $t\in[s_{1},s_{2}]$, $0\leq s_{1}<s_{2}\leq T$ and let $\mathbf{X}_{[s_{1},s_{2}]}:=\{X_{t}\}_{t\in[s_{1},s_{2}]}$, and $\mathbf{W}_{[s_{1},s_{2}]}:=\{W_{t}\}_{t\in[s_{1},s_{2}]}$. We will use a measure disintegration of the diffusion path measure with respect to a sampled value at time $s_{2}$. Suppose $f_{t,s_{2}}(x^{\prime}|x)$ denotes the unknown transition density from time $t$ to $s_{2}$ associated to (4.1). An equivalent formulation for the law of $\mathbf{X}_{[s_{1},s_{2}]}$ would be to sample from $f_{s_{1},s_{2}}$ to obtain $(x,x^{\prime})\in\mathbb{R}^{2d_{x}}$ and then to interpolate between $x$ and $x^{\prime}$ or "fill in" the path using an alternative diffusion that is conditioned to arrive at point $x^{\prime}$ at time $s_{2}$. We will refer to this this diffusion process as the bridge process and its dynamics will be will be as in (4.1), but instead its drift will be given by $\mu(x)+\Sigma(x)\nabla_{x}\log{f}_{t,s_{2}}(x^{\prime}|x)$. Let $\mathbb{P}_{x,x^{\prime}}$ denote the law of the solution of the SDE (4.1), on $s_{1},s_{2}]$, started at $x$ and conditioned to hit $x^{\prime}$ at time $s_{2}$.
@@
(4.2) is specified so that its transition density $\tilde{f}$ is available; see [24, Section 2.2] for the technical conditions on $\tilde{\mu},\tilde{\Sigma},\tilde{f}$. The main purpose of $\{\tilde{X}_{t}\}_{t\in[s_{1},s_{2}]}$ is to sample $x^{\prime}$ and use its transition density to construct another process $\{X_{t}^{\circ}\}_{t\in[s_{1},s_{2}]}$ conditioned to hit $x^{\prime}$ at $t=s_{2}$; which in turn will be an importance proposal for $\{X_{t}\}_{t\in[s_{1},s_{2}]}$. Set:
@@
The authors in [24] prove that $\mathbb{P}_{x,x^{\prime}}$ and $\mathbb{P}^{\circ}_{x,x^{\prime}}$ are absolutely continuous w.r.t. each other, with Radon-Nikodym derivative:
$\frac{d\mathbb{P}_{x,x^{\prime}}}{d\mathbb{P}^{\circ}_{x,x^{\prime}}}(\mathbf{X}_{[s_{1},s_{2}]})=\exp\Big\{\int_{s_{1}}^{s_{2}}L_{s_{2}}(t,X_{t})dt\Big\}\frac{\tilde{f}_{s_{1},s_{2}}(x^{\prime}|x)}{f_{s_{1},s_{2}}(x^{\prime}|x)},$ (4.5)
@@
with $\textrm{Tr}(\cdot)$ denoting the trace of a squared matrix.
We also need to include the sampling of $x^{\prime}$, so set
$R_{s_{1},s_{2}}(\mathbf{X}_{[s_{1},s_{2}]}):=\frac{d\mathbb{P}_{x,x^{\prime}}}{d\mathbb{P}^{\circ}_{x,x^{\prime}}}(\mathbf{X}_{[s_{1},s_{2}]})f_{s_{1},s_{2}}(x^{\prime}|x),$
where one should note that remarkably $R_{s_{1},s_{2}}$ does not depend on the unknown transition density $f_{s_{1},s_{2}}$.